OpenStudy (anonymous):

a runner runs around a circular track with a circumference of 1000 meters . find the distance covered and the displacement made by the runner ?

OpenStudy (mathstudent55):

If the runner runs 1 lap, then the distance covered is 1000 meters. The displacement is 0 because he is back where he started.

OpenStudy (anonymous):

ammm, is there any solution ? like formulas ?

OpenStudy (anonymous):

thank you so much po for replying :)

OpenStudy (mathstudent55):

You didn't say how many laps he runs.

OpenStudy (anonymous):

that was given by our teacher po :)

OpenStudy (anonymous):

Im gonna aswer whats your answer Mr, or Ms. :) Thank you so much , :)

OpenStudy (mathstudent55):

Ok. If he runs 1 lap, you have the distance covered and the displacement above. What if he runs half a lap? Then the distance covered is half of the circumference, or 500 meters. The displacement is the diameter of the circle. Then you can calculate the diameter using the circumference. \(C = 2\pi r\) \(C = \pi d\) \(d = \dfrac{C}{\pi} \) \(d = \dfrac{1000}{\pi} \)

OpenStudy (anonymous):

wow.:) thanks po for your time :) can i ask again po?

OpenStudy (mathstudent55):

I have to go now. I'll be back later in the day.

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (anonymous):

okay , thank you po :) thank you so much :)